Identify the type of equations and Solve

a) 𝑦′ = π‘₯𝑦 1+π‘₯2

b) sin 𝑦 𝑑π‘₯ + cos 𝑦 𝑑𝑦 = 0

c) 𝑑𝑦 = 2𝑑 (𝑦2 + 9 )𝑑𝑑


1
Expert's answer
2021-05-28T06:10:05-0400

"a)\newline
 𝑦′ = \frac{𝑥𝑦}{1+𝑥^2}\newline
\text{This linear ordinary Differential equation.}\newline\text{
Solving the given differential equation by variable separable form,}\newline
 \frac{dy}{y}= \frac{𝑥}{1+𝑥^2}dx\newline
\text{
Integrate both side,}\newline
\int\frac{dy}{y}= \frac{1}{2}\int\frac{2𝑥}{1+𝑥^2}dx\newline
lny= \frac{1}{2}ln(1+x^2)+c\newline


b)\newline
 sin 𝑦 𝑑𝑥 + cos 𝑦 𝑑𝑦 = 0\newline\text{
This linear ordinary Differential equation}.\newline\text{
Solving the given differential equation by variable separable form,}\newline
 \frac{cosy}{siny}dy= -dx\newline\text{
Integrate both side,}\newline
\int\frac{cosy}{siny}dy= -\int dx\newline
lnsiny=-x+c
\newline
c)\newline
 𝑑𝑦 = 2𝑡 (𝑦^2 + 9 )𝑑𝑡\newline\text{
This linear ordinary Differential equation}.\newline\text{
Solving the given differential equation by variable separable form,}\newline
𝑑𝑦 = 2𝑡 (𝑦^2 + 9 )𝑑𝑡\newline
 \frac{1}{3^2+y^2}dy= 2tdt\newline
\frac{1}{9} \frac{1}{1+(\frac{y}{3})^2}dy= 2tdt\newline\text{
Integrate both side,}\newline
\frac{1}{9}\int \frac{1}{1+(\frac{y}{3})^2}dy= 2\int tdt\newline
\frac{1}{3} tan^{-1}(\frac{y}{3})= t^2+c\newline"